OpenStudy (anonymous):

y^2+3yz+8z-4x=0 solve for z?

OpenStudy (anonymous):

If the last one 'made sense' why are you confused about how to do this one?

OpenStudy (anonymous):

beacuse i'm a slow learner and this one is different. sorry if i'm bad at math. that's kinda why i'm asking for help

OpenStudy (anonymous):

This one isn't any different.

OpenStudy (anonymous):

islote all the terms that have a z factor in them.

OpenStudy (anonymous):

Isolate rather

OpenStudy (anonymous):

yes it is! there's exponents

OpenStudy (anonymous):

None of the exponents are on the z..

OpenStudy (anonymous):

so it'd be y^2-4z=3yz-8z?

OpenStudy (anonymous):

Ok, do you know what I mean by isolate the terms with z factors?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok. Do you know what I mean by 'term'?

OpenStudy (anonymous):

Yes, I'm not *that* oblivious

OpenStudy (anonymous):

Ok so which terms don't have z factors?

OpenStudy (anonymous):

I don't get what you mean by isolate them, I thought that's what I did

OpenStudy (anonymous):

I'm not meaning to be insulting. I just need to know what you know and what I should explain

OpenStudy (anonymous):

No no I understand, I'm just frustrated is all So the z terms are 3yz and the 8z

OpenStudy (anonymous):

Yes. Good.

OpenStudy (anonymous):

Now to isolate them means to subtract (or add) all the other terms that don't have a z to the other side of the equal sign.

OpenStudy (anonymous):

Oh, you actually did that before but I got confused because you said -4z instead of -4x

OpenStudy (anonymous):

I'm sorry. You were right so far: \[y^2 - 4x = -3yz - 8z\]

OpenStudy (anonymous):

Oh whoops sorry, mistype. That makes more sense then cuz I was confused cuz I thought I did it right lol

OpenStudy (anonymous):

Now. The next step is to factor out the z from each of the terms that have it.

OpenStudy (anonymous):

So instead of -3yz - 8z we have z(-3y - 8)

OpenStudy (anonymous):

Okay got that

OpenStudy (anonymous):

we divide the z away from each term and instead put it out in front

OpenStudy (anonymous):

because you can see that \[z(-3y - 8 ) = -3yz - 8z\] So we are just going in reverse

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Now to get the z by itself we just divide both sides of the equal sign by the non-z factor

OpenStudy (anonymous):

So it'd be z(3y-8)/-4-y^2? Or am I totally lost?

OpenStudy (anonymous):

Right track. Wrong factor

OpenStudy (anonymous):

Instead divide both sides by (-3y-8)

OpenStudy (anonymous):

the factor next to the z we want to get rid of, so we divide it off, but we have to divide both sides to keep it equal

OpenStudy (anonymous):

Oh because their the terms you factored the z from? Got it

OpenStudy (anonymous):

\(y^2 - 4x = z(-3y - 8)\)\[\implies \frac{y^2 - 4x}{-3y - 8}= z\frac{-3y-8}{-3y-8}\]

OpenStudy (anonymous):

and the fraction on the left has the same value on top and bottom, so that is just 1 now.

OpenStudy (anonymous):

err fraction on the right, sorry

OpenStudy (anonymous):

\[\implies \frac{y^2 - 4x}{-3y - 8}= z \]

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

Yes I think so

OpenStudy (anonymous):

Could you help me with more of my packet?? You're explaining better than any of my teachers ever have. If you don't have time, that's totally fine tho

OpenStudy (anonymous):

Ok so here's a summary of what we did. You should write this down in your notes. Then see if you can do it for yourself the next time. Solving equation for 'z' (could be any letter) 1) Isolate the terms that have a factor of z by adding/subtracting the z terms to one side and the non-z terms to the other side. 2) factor out the z from each term, putting it in front of parens and dividing the z off each term 3) divide both sides by the factors being multiplied by your z

OpenStudy (anonymous):

I'm happy to help

OpenStudy (anonymous):

feel free to ask more questions

OpenStudy (anonymous):

Wonderful! Thank you. I'm new to this site, so do you want the q's separately or just on here?

OpenStudy (anonymous):

separately is better otherwise it gets lagged when the thread gets too long